OpenStudy (anonymous):

how do you know when a polynomial is completely factored? like... 4(n^4 -8) or 5x^4 (12x^3 - 6x^2 - 18x)

OpenStudy (anonymous):

A polynomial is completely factored when you can't do any more transformation . 5x^4 (12x^3 - 6x^2 - 18x). Put 6x in factor. 30 x^5 (2x^2 - x - 3). Put the trinomial into factors use Tip 1 when you have a + b + c = 0 10x^5 *(x - 1)(2x + 3). This is completely factored. 4(n^4 - 8) = 4 (n^2 - 2V2)(n^2 + 2V2).